### v2

# When Collaborative Maintenance Falls Short: The Persistence of Retracted Papers on Wikipedia

Abstract
Wikipedia serves as a key infrastructure for public access to scientific knowledge, but it faces challenges in maintaining the credibility of cited sources—especially when scientific papers are retracted. This paper investigates how citations to retracted research are handled on English Wikipedia. We construct a novel dataset that integrates Wikipedia revision histories with metadata from Retraction Watch, Crossref, Altmetric, and OpenAlex, identifying 1,181 citations of retracted papers. We find that 71.6% of all citations analyzed are problematic. These are citations added before a paper’s retraction, as well as the citations introduced after retraction without any in-text mention of the paper’s retracted status. Our analysis reveals that these citations persist for a median of over 3.68 years (1,344 days). Through survival analysis, we find that signals of human attention are associated with a faster correction process. Unfortunately, a paper’s established scholarly authority—a higher academic citation count—is associated with a slower time to correction. Our findings highlight how the Wikipedia community supports collaborative maintenance but leaves gaps in citation-level repair. We contribute to CSCW research by advancing our understanding of this sociotechnical vulnerability, which takes the form of a community coordination challenge, and by offering design directions to support citation credibility at scale.
Keywords:
collaborative maintenance, Wikipedia, retracted papers, sociotechnical systems, knowledge infrastructure, scientific misinformation
1 Introduction
Wikipedia is one of the most influential collaborative knowledge infrastructures in the world. Maintained by a decentralized network of human contributors, automated bots, and community governance processes, it produces and maintains a vast, continuously evolving corpus of public knowledge forte2013defining ; halfaker2012bots ; geiger2010work . This content shapes public understanding across scientific, cultural, and political domains. It is increasingly used not only by students and journalists but also by search engines and large language models as a foundational information source touvron2023llama ; brown2020language . Given its centrality in the modern knowledge ecosystem, the ability of the Wikipedia community to maintain the integrity of its information over time has profound epistemic and societal consequences fallis2008toward ; aragon2021preliminary .
A growing body of CSCW and HCI research has examined how the Wikipedia community responds to real-world change, particularly in the context of high-attention events such as natural disasters, political controversies, social movements, and celebrities’ deaths keegan2013hot ; keegan2012staying ; twyman2017blm ; keegan2015is . These studies have shown that Wikipedia can mobilize contributors to rapidly produce and update information in response to new developments, demonstrating the community’s capacity for collaborative sensemaking. However, far less is known about how well Wikipedia’s editors and community processes handle low-salience epistemic disruptions like paper retractions—cases where previously accepted knowledge becomes invalidated after its initial citation on the platform. These disruptions are slower, less visible, and often go unnoticed by editors and readers alike. The persistence of such trackable errors demonstrates a fundamental coordination challenge within large-scale, decentralized communities: how to ensure that critical updates are routed to the specific locations where repair is needed kittur2008harnessing ; forte2012coordination .
Figure 1: Estimated cumulative and monthly page views of English Wikipedia articles that cite retracted papers without in-text mention of retraction status. The figure only shows citations with no accompanying retraction-related keywords (“retract”, “withdraw”, or “invalid”) in the main text. This keyword-based detection of uncorrected retraction provides a coarser approximation than the correction categorization used in the main analysis.
Our study focuses on scientific retractions, one of the clearest signals of post-publication epistemic breakdown jackson201411 ; fallis2008toward . Retractions occur when scientific findings are withdrawn due to error, misconduct, or other credibility issues, and they function as formal invalidations of prior knowledge fang2012misconduct . When such papers remain cited on Wikipedia without mention of retraction status or removal, they present a critical test of the community’s ability to maintain trustworthy knowledge. From the perspective of readers and other downstream users, this poses serious risks. A person seeking information about a medical treatment, scientific finding, or public health claim may unknowingly encounter content on Wikipedia grounded in invalidated evidence, without any clear warning or correction. Such lapses not only risk misinforming individuals in critical decision-making contexts but also erode trust in Wikipedia as a reliable source of up-to-date knowledge.
As is illustrated in Figure 1, English Wikipedia articles containing uncorrected citations to retracted papers have received over 250 million cumulative page views as of 10/24/2024. While not every view results in exposure to the specific citation or misinformation, this figure reflects the substantial reach of flawed content. Moreover, given Wikipedia’s role as a foundational information layer across the web, including in search engine results, educational settings, and machine learning training data, the consequences of uncorrected retractions extend far beyond individual pages and would be difficult to quantify thelwall2025does .
Rather than analyzing editor behavior directly, this paper takes a system-level perspective on how the Wikipedia community handles retracted scientific knowledge. Building on CSCW scholarship on collaborative infrastructures, knowledge governance, and epistemic maintenance ackerman2013sharing ; aragon2021preliminary ; forte2012coordination , we examine how citations of retracted papers are distributed across the platform and what observable features are associated with how quickly these citations are corrected. To do so, our study focuses on corrective actions that are directly visible to a typical reader within the main body of an article. We therefore define a “correction” narrowly as either the complete removal of the citation or the addition of a visible in-text warning, which includes explicit mentions of the retraction or editorial notes indicating the need for better sources. This definition intentionally excludes automated templates added only to the reference section, which may not be visible to readers who do not consult the bibliography. This distinction is crucial because it recognizes the division of labor within the Wikipedia community. Automated tools are essential for the scalable work of detection and flagging halfaker2013rise ; halfaker2012bots , providing a vital signal primarily for an editorial audience. However, a complete, reader-facing repair often requires a subsequent act of contextual judgment and editing by humans. Since this human oversight is a key variable and not always guaranteed kharazian2024governance ; baigutanova2023longitudinal , our study’s focus on in-text corrections is designed to measure the success of this complete human-in-the-loop process.
Here, we combine data from Retraction Watch oransky2012retraction , Crossref Events hendricks2020crossref , Altmetric LLC altmetric , OpenAlex priem2022openalex , and Wikipedia revision and page view histories to investigate two core aspects of retracted research on Wikipedia:
  • Overview of retracted paper citations on the English Wikipedia. We explore the distribution of uncorrected citations to retracted papers relative to non-retracted ones and identify the academic domains most susceptible to citing retracted papers on Wikipedia. This descriptive overview quantifies the scope of the issue and its potential impact on English Wikipedia users.
  • What factors influence the speed at which retracted paper citations on English Wikipedia are corrected? Our study investigates the characteristics associated with citations of retracted papers being corrected on Wikipedia. We analyze various features, including a range of features related to the paper itself (e.g., its academic domain, open access status, and academic citation count). Additionally, we examine factors related to the retraction event (e.g. the reasons for retraction, time from publication to retraction, and editorial events before retraction). Finally, we analyzed Wikipedia-specific factors such as the level of editorial activity (e.g., the number of unique editors). These factors provide insight into the challenges of identifying and addressing problematic citations.
Our findings show that Wikipedia’s responsiveness to citations of retracted papers is inconsistent and often slow. While some citations are corrected quickly, many persist uncorrected for years. Among the 1,181 citation pairs analyzed, we found 71.6% to be problematic—either added before the retraction or introduced after the retraction without in-text retraction warnings. Citations are corrected more quickly when the retraction was for serious issues like fraud or error, when Wikipedia editors flagged the paper as controversial at the time of its initial citation, when the paper is open access, and when it received substantial attention on Twitter/X. Corrections are also faster on pages with more unique editors and for papers in the Physical and Social Sciences. Conversely, citations to papers with higher academic citation counts are corrected more slowly, suggesting that a paper’s established impact may create editorial inertia. These patterns demonstrate the complex interplay of visibility, editorial capacity, and contextual salience in Wikipedia’s citation maintenance practices.
Our paper contributes to CSCW research on collaborative knowledge infrastructures in four ways. First, it shifts the focus from Wikipedia’s well-documented responsiveness to high-salience events to its underexamined handling of low-visibility epistemic breakdowns, such as retracted scientific papers. Second, it provides a large-scale empirical analysis of how retracted citations persist—or are corrected—across academic domains, revealing how features such as social media signals, retraction reasons, and editor numbers are associated with repair outcomes. Third, it highlights a disconnect between reference-level flagging and substantive in-text repair. By focusing our analysis on corrections visible within the article body, we highlight how automated signals in the reference section may not translate into the human editorial work required to address problematic citations for a general reader appropriately. This indicates a broader design gap between machine-generated alerts and editorial workflows. Finally, it provides actionable opportunities for design, suggesting how Wikipedia’s tools, workflows, or interfaces might better support editors in identifying and repairing citation-level credibility failures.
2 Related Work
2.1 Wikipedia as an Epistemic Infrastructure
Wikipedia has served as a prime example in CSCW and HCI research for studying large-scale, decentralized collaboration. As a social computing system, it depends on the interplay of human contributors, automated bots, and governance policies to produce and maintain public knowledge forte2013defining ; halfaker2012bots ; geiger2010work . This sociotechnical infrastructure supports a complex ecosystem of roles, routines, and technical tools that enable collaborative knowledge production ackerman2013sharing ; forte2012coordination .
Prior work has demonstrated how Wikipedia adapts to rapidly changing external events. In crisis contexts such as natural disasters or breaking news, editors engage in high-tempo coordination, rapidly creating and revising articles to reflect emerging realities keegan2013hot ; keegan2012staying . Studies of social movement coverage, such as the documentation of Black Lives Matter, extend this perspective to the formation of collective memory in contentious and evolving domains twyman2017blm . Keegan et al. keegan2015is show how editors update Wikipedia biographies after someone dies, working together to correct facts while also honoring the person’s memory. Together, these studies demonstrate Wikipedia’s potential for responsiveness, focusing on high-salience events that trigger rapid editorial attention.
Yet recent work highlights Wikipedia’s vulnerabilities in less visible contexts. Kharazian et al. kharazian2024governance show how, in smaller language editions of Wikipedia, editorial control can become concentrated in the hands of a few users, allowing misinformation to persist. Similarly, Aragón et al. aragon2021preliminary identify risks to knowledge integrity that can arise even without deliberate manipulation. Together, these studies show that Wikipedia’s openness—while foundational to its success—also means that no individual contributor is accountable for the accuracy of any specific claim. When responsibility is diffuse and editorial attention is uneven, epistemic breakdowns can persist without being noticed or repaired. Moreover, Wikipedia’s information credibility can face intentional challenges. Kumar et al. kumar2016disinformation studied the creation and detection of Wikipedia hoaxes, fabricated articles designed to deceive readers. Our study contributes to the literature by examining how Wikipedia handles a particular form of non-adversarial epistemic disruption: the citation of scientific articles that have been formally retracted. Failure to maintain the credibility of cited scientific sources leads to a non-negligible breakdown in information quality that is impactful regardless of intention.
Prior work has provided important insights into Wikipedia’s collaborative maintenance processes at the article level. However, fewer studies have examined how the platform manages the credibility of cited knowledge when it becomes formally invalidated. Rather than tracing editorial workflows directly, we analyze the distribution of these citations across Wikipedia and identify factors associated with whether they are eventually corrected or persist unaddressed. This approach highlights how certain infrastructural and visibility features—such as the number of unique human editors, Twitter/X attention to the papers, and the papers’ academic domains—shape citation repair in decentralized collaborative systems.
2.2 Scientific Knowledge and Citations on Wikipedia
Scientific knowledge plays a central role in Wikipedia’s coverage of specialized topics, and recent scholarship has explored how scientific sources are cited, structured, and maintained. Prior work has mapped the landscape of scientific citations on Wikipedia, revealing both the breadth of disciplines represented and the dominant role of biomedical research and papers from high-impact journals. Singh, West, and Colavizza singh2021wikipedia quantified this trend, finding that the life sciences and biomedicine dominate the cited literature, with journals such as Nature and Science among the most frequently referenced. Arroyo-Machado et al. arroyo2020science similarly found that multidisciplinary and biomedical journals dominate Wikipedia’s scientific references. Nicholson et al. nicholson2021measuring assessed citation quality, showing that Wikipedia tends to cite reputable journals, but still includes a notable fraction of sources with limited academic validation. Other studies have documented representational biases, such as the under-citation of women and non-Anglophone authors zheng2023gender .
Complementing this citation-level research, scholars have also examined how Wikipedia’s article quality ratings relate to citation behavior. Arroyo-Machado et al. arroyo2022wikinformetrics found that articles rated as higher quality by the Wikipedia community tend to include more academic references, attract more edits, and receive higher page views—but that these associations vary depending on the topical domain. For example, the authors found that a high number of academic references strongly correlates with scholarly-oriented articles in fields like medicine and history. In contrast, high page views are more typical of articles about popular culture, and high talk page activity is associated with controversial subjects. More recently, Baigutanova et al. baigutanova2023longitudinal conducted a large-scale longitudinal study of Wikipedia’s reference quality, finding that sourcing practices have generally improved over time, particularly with increasing reliance on peer-reviewed journals and scientific literature. This trend toward higher-quality sourcing builds on earlier observations about Wikipedia’s increasing preference for high-impact journals such as Nature and Science nielsen2007scientific ; wedemeyer2008quality . Nevertheless, concerns persist about the reliability and maintenance of references, particularly in articles that receive limited editorial attention.
While prior research has primarily focused on the inclusion and quality of citations, relatively little is known about how Wikipedia handles erroneous scientific citations, particularly those referencing retracted research. The study most closely aligned with ours is by Nicholson et al. nicholson2021measuring , who identified 118 retracted articles cited across 297 Wikipedia pages. In a manual sample of 50 citations, they found that about half failed to acknowledge the retraction, and even when retractions were mentioned, they were typically used illustratively rather than corrected or flagged clearly. Additionally, their study was based on a small sample. It did not examine what factors drive citation correction, thus creating an opportunity for a large-scale characterization of retraction persistence patterns across Wikipedia.
We build on this line of work by identifying retracted papers cited on Wikipedia and analyzing how those citations are handled over time through article revisions. Our study offers a large-scale account of how retracted citations are noted, ignored, or removed. We also model the factors that shape these outcomes. In doing so, we extend prior research on citation practices by shifting attention from how scientific citations are sourced and evaluated to how they become invalid and are subsequently addressed. Our work also contributes to CSCW scholarship on collaborative maintenance and infrastructure resilience by examining how Wikipedia editors manage epistemic failures at the citation level.
2.3 Retractions, Online Attention, and Challenges of Correction
Retractions are intended to correct the scientific record, but their corrective effects are often limited once flawed research has circulated online. Prior work shows that retracted papers frequently receive more online attention compared to their non-retracted counterparts, particularly on social media, news platforms, and in knowledge repositories peng2022dynamics . This attention is often exhausted before the retraction is issued, and, in some cases, the retraction itself may unintentionally amplify visibility. While increased attention can surface public criticism—especially on platforms like Twitter/X peng2022dynamics ; amiri2024tweeters —only a small and selective group of users tends to engage in such critical discussions. Moreover, retracted papers have also been shown to disseminate false information, particularly around sensitive topics such as vaccination, thereby undermining public trust in science abhari2024they . Emerging work has explored whether language models can help flag problematic papers based on these online signals, with promising results using Twitter/X data zheng2024can .
While these studies illuminate how retractions garner online visibility, they primarily focus on digital and social media. Much less is known about how collaborative knowledge systems like Wikipedia respond to retracted science. We address this gap by examining retraction correction as a form of sociotechnical repair, and our work contributes to CSCW by extending its focus on knowledge governance, the breakdown of citation credibility, and maintenance in decentralized systems.
3 Empirical Setup
3.1 Data Collection
Figure 2: Overview of the data collection and annotation process. The left panel shows the construction of the main dataset, including the extraction of 1,181 retracted paper–Wikipedia citation pairs from Crossref Event Data and corresponding revision histories from Wikipedia. Manual annotation determined whether each citation was cited with its retraction status noted at its first appearance and corrected in the latest revision. The right panel displays the separate dataset of all Wikipedia-cited papers (N = 1,590,982), which is used to calculate citation ratios across academic domains. Retraction status is linked using Retraction Watch, and domains are categorized via OpenAlex.
We integrated data from Retraction Watch oransky2012retraction , Crossref Events hendricks2020crossref , Altmetric LLC altmetric , OpenAlex priem2022openalex , as well as Wikipedia’s page view statistics (Wikimedia REST API) and revision histories (MediaWiki Action API). This integration enables us to systematically examine how retracted research is cited and handled on Wikipedia, allowing us to investigate further how features of papers, Wikipedia pages, online mentions, and temporal trends (e.g., whether citations occurred before or after retraction, and the time between publication and retraction) are associated with the correction of retracted paper citations.
The Retraction Watch database is the most comprehensive and widely adopted resource on retracted articles brainard2018rethinking , indexing all retracted Digital Object Identifiers (DOIs) and their metadata, such as retraction date, publication date, and retraction reason. In this study, we used Retraction Watch to obtain a list of retracted papers and their associated metadata, including article titles, retraction dates, publication dates, and retraction reasons. To ensure our analysis focused exclusively on retracted papers, we excluded the ones with expressions of concern or corrections. Expressions of concern serve as warnings while investigations are ongoing, and corrections address minor issues without invalidating the entire study. Retractions, by contrast, formally withdraw a paper from the academic record.
We used the Crossref Event API to collect all Wikipedia pages that have cited the retracted DOIs. Crossref periodically crawls Wikipedia and logs pages that contain outbound links to Crossref DOIs, along with detection timestamps and metadata hendricks2020crossref . The Crossref Event API is the most accurate in indexing Wikipedia citation data ortega2018reliability , and only indexes papers with valid DOIs. 833 retracted papers were cited on 900 Wikipedia pages, forming 1,181 retracted paper–Wikipedia page citation pairs. To accurately date when the retracted DOI was introduced to the Wikipedia pages and when the retracted DOI was properly corrected on Wikipedia (complete removal or adding an in-text warning), we used the MediaWiki Action API to retrieve and analyse all revisions ($N=824,012$) of these pages. This allowed us to perform a detailed temporal analysis of how retracted papers were cited and subsequently addressed on Wikipedia. To ensure our analysis was relevant to public-facing content, we excluded pages from non-article namespaces such as Talk, Sandbox, Draft, and User.
We constructed a dataset separate from the one used in our main analysis to compare the academic domain distribution of retracted and non-retracted papers cited on Wikipedia. For consistency, we identified 538 retracted papers using the latest Altmetric data dump (current as of August 14, 2023), rather than relying on our original Wikipedia retracted citations dataset. This allowed us to align data sources when comparing them to the 1,590,444 non-retracted papers also cited on Wikipedia and indexed in Altmetric. The Crossref Event API proved impractical for extracting such a large volume of citations to non-retracted papers, so relying on Altmetric for both retracted and non-retracted papers ensured source consistency for this descriptive analysis. This dataset is used solely for comparing academic domain distributions and is not part of our modeling of correction outcomes. To enrich the dataset, we retrieved academic domain classifications from OpenAlex, along with additional metadata such as open access status, citation counts, and co-author team sizes. All data were collected in October 2024.
3.2 Types of Retracted Paper Citations
When retracted papers are first introduced as citations on Wikipedia, they can appear in the following ways:
  • Cited Before Retraction. These are instances where papers are cited on Wikipedia before their official retraction date. At the time of citation, the paper’s retracted status was not yet known, and no corrective context was provided in the article. If these citations remain after retraction without being updated or flagged, they risk misleading readers.
  • Cited After Retraction Without In-Text Acknowledgment. These citations occur after the paper has been officially retracted, but there is no indication of the retraction in the Wikipedia article. Readers may mistakenly assume that the paper’s results are valid, which could lead to misconceptions.
  • Cited After Retraction With In-Text Acknowledgment. In some cases, retracted papers are introduced with explicit acknowledgment of their retraction status. This definition intentionally excludes automated warnings in the reference section, as our focus is on the human editorial actions that make a retraction’s status transparent to a reader of the main article text.
For this study, we consider problematic citations to be the ones Cited Before Retraction and Cited After Retraction Without In-Text Acknowledgment. These scenarios pose significant risks of misinformation, as readers may unknowingly rely on invalid research. We focus exclusively on whether retraction status is clearly mentioned in the article text. Merely critiquing a paper or casting doubt on its findings does not meet our threshold for acknowledgment unless the retraction is explicitly stated.
To identify these cases, we tracked when each retracted paper first appeared in its associated Wikipedia articles by searching for its DOI and title across revision histories. One of the authors then manually examined whether the retraction status was mentioned in the main text of those initial citing revisions.
3.3 Handling Problematic Retracted Paper Citations
By manually inspecting the latest version of pages containing problematic citations in our dataset, we observe three ways of handling problematic retracted paper citations on Wikipedia:
  • Complete Removal. The most straightforward approach to correcting citations of retracted papers has been to remove them entirely from the Wikipedia article. This method eliminates any potential for future misinformation stemming from the problematic source. However, it can be controversial if the retracted paper contributed historically significant or contextually relevant information to the article. In academic paper writing, the complete removal of retracted paper citations is always chosen when editors and reviewers do not consider its continued presence essential retractionwatch2018how .
  • In-Text Retraction Acknowledgment. A second approach involves placing a clear warning near the citation in the Wikipedia article text, ideally in the same paragraph where the retracted paper is referenced. This method informs readers about the paper’s retracted status while preserving the citation for transparency or historical context. Such warnings may appear as editorial notes or parenthetical remarks, e.g., “(This paper has been retracted).” This category also includes other explicit in-text editorial flags, such as the better source needed template. This strategy balances accuracy and informational completeness, but it relies heavily on editors’ consistency and awareness.
  • Reference Section Acknowledgment Only. In many cases, a paper’s retracted status is only acknowledged in the reference section, without modifying the article’s main text. Because these reference-level warnings are less visible to casual readers—who may not consult the references—this practice allows potentially invalidated claims to persist unchallenged in the main body of the article. As a result, critical misinformation can remain embedded in Wikipedia’s content. For example, automated tools like RetractionBot Headbomb2024Signpost are designed to flag retracted papers on Wikipedia by adding annotations primarily to the reference section.
For our analysis, we consider citations to be corrected if it was Complete Removal or In-Text Retraction Acknowledgment, as these approaches offer high visibility to readers.
To identify these two types of corrections, we employed a two-step process. First, to find instances of Complete Removal, we searched the revision history of each article for the titles and DOIs of the retracted papers. Second, for all remaining cases where the citation was still present, we manually reviewed the most recent revision to identify instances of In-Text Retraction Acknowledgment, assessing whether the retraction status was explicitly mentioned in the main text.
3.4 Analytical Approach: Survival Analysis
Our analysis seeks to understand the factors that influence how quickly problematic citations are corrected. A significant portion of these citations remained uncorrected at the end of our data collection period. This type of incomplete observation is known as right-censored data. Standard regression models are not suited to handle such data. Therefore, to properly account for these censored observations, we employ survival analysis, a method well-suited for modeling time-to-event data in online communities ma2017write ; mindel2025timely . This method enables us to model the time until a correction occurs (the “event”) while appropriately incorporating information from both corrected and uncorrected (censored) citations.
3.5 Dependent Variable: Time to Correction
Our survival model analyzes the time it takes for a problematic citation to be corrected on Wikipedia. The primary dependent variable is the duration, measured in days, from when a citation becomes “at risk” of correction-that is, after it is both present on Wikipedia and officially retracted—until the correction event occurs.
Crucially, the start of the risk period for each citation is defined as the later of two dates: the paper’s official retraction date or the date the citation was first added to the Wikipedia page. This ensures that a citation is only considered “at risk” after it is both present on Wikipedia and officially retracted. As a data cleaning step, we excluded a small number of “proactive corrections”—cases where a correction was made before the official retraction date. To avoid numerical issues in parametric survival fitting, all durations were adjusted by adding a small $\epsilon=10^{-3}$ days.
A correction event is defined as either the complete and persistent removal of the citation or the addition of an in-text warning of the retraction in the immediate context of the citation. For citations that remained uncorrected as of our data collection endpoint (October 24, 2024), they are treated as right-censored observations. Their duration is the time from the start of their risk period until this endpoint.
3.5.1 Correction via In-Text Retraction Acknowledgment
For citations that were eventually corrected by mentioning the retraction status in the main text of the article, we begin by identifying a retraction-related keyword in the final, corrected revision of the Wikipedia article. These may include general terms like “retracted” or “withdrawn,” or more specific phrases such as “The article was retracted on 13 August 2010.” Using this final revision, we select a representative keyword or phrase associated with the correction and trace backward through the revision history to identify the earliest revision where that keyword appeared.
To avoid mistaking short-term rewording or editing noise as correction reversals, we required that the identified correction keyword or phrase, once added, remain in subsequent revisions, allowing for a temporary disappearance of up to three consecutive revisions. This heuristic balances robustness against transient noise with sensitivity to meaningful editorial behavior. Because articles may cite multiple retracted papers, we manually verify that the keyword refers to the specific citation of interest. If needed, we extract a more precise phrase from the later revision that unambiguously refers to the retracted paper in question.
3.5.2 Correction via Removal
For citations corrected through removal, we apply a similar logic. A citation is considered removed when it is absent from both the article’s main text and its reference section. We identify the earliest revision in which the citation is no longer present and remains removed, again defined as allowing for up to three subsequent revisions where the citation might briefly reappear.
3.5.3 Assumptions
Our approach to identifying correction events, which is based on a dual-track search for either Complete Removal or In-Text Retraction Acknowledgment, relies on several working assumptions:
  • Wikipedia editors typically choose either in-text retraction acknowledgment or removal as a correction strategy for retracted citations.
  • Once corrected, a citation remains corrected in that same form throughout subsequent revisions.
  • The retraction-related keyword or phrase, once added, remains consistent in subsequent revisions.
  • The presence or absence of a citation is reliably reflected in the article’s main text and references.
By applying this methodology, we estimate the time from retraction to correction for each citation, enabling us to assess both the speed and consistency of Wikipedia’s response to retracted science.
3.6 Independent Variables
Table 1: Summary statistics of independent variables used in the survival analysis. Variables marked with “(at risk start)” were measured at the beginning of each citation’s unique risk period, which starts at the later of two dates: the paper’s official retraction date or the date the citation was added to Wikipedia.
Mean ($\mu$) Std Dev ($\sigma$) Min Max Median
Paper Features
   Open Access 0.62 0.48 0 1 1
   Academic Citation Count (at risk start) 147.77 417.75 0 5436 46.00
   Team Size 6.34 4.65 0 28 5.00
   Physical Sciences (domain) 0.13 0.34 0 1 0
   Life Sciences (domain) 0.36 0.48 0 1 0
   Social Sciences (domain) 0.08 0.27 0 1 0
   Tweet Mention Count (at risk start) 345.26 2159.96 0 33946 2.00
Retraction Event Features
   Time to Retraction (days) 2409.45 2096.41 0 11344 1914.50
   Pre-Retraction Editorial Event 0.12 0.33 0 1 0
   Retraction Reason (salient) 0.68 0.74 0 2 1.00
Wikipedia Page Features
   Number of Unique Editors (at risk start) 151.79 275.27 0 2343 34.50
   Controversy Noted at Initial Citation 0.05 0.22 0 1 0
To identify factors that accelerate or decelerate the correction of retracted paper citations on Wikipedia, we examine four different groups of characteristics, as follows (see Table 1):
3.6.1 Paper Features
Paper characteristics are particularly important because they capture the intrinsic properties of the publications that may either facilitate or hinder their detection and correction. These data were retrieved from external scholarly databases, including OpenAlex priem2022openalex , Retraction Watch oransky2012retraction , and Altmetric altmetric .
  • Open Access. Open access status is coded as a binary variable. This variable allows us to examine whether open-access availability affects the speed with which retracted papers are corrected on Wikipedia. Prior research suggests that open-access papers may receive greater visibility and more citations due to their unrestricted availability piwowar2018state ; teplitskiy2017amplifying . Zheng et al. found that open access may help speed up the detection and retraction of flawed articles in biochemistry zheng2024can .
  • Academic Citation Count. Academic citations serve as a proxy for a paper’s scholarly impact and visibility. To avoid look-ahead bias, we measured this variable dynamically for each citation pair. Specifically, we only counted citations the paper had accumulated before the start of its risk period on Wikipedia, ensuring our measure of scholarly impact could not be influenced by the correction process itself. Using the OpenAlex API, we determined the number of citations a retracted paper had received up to the start of each citation’s unique risk period (i.e., the point after it was both retracted and cited on Wikipedia). While a paper’s higher citation count could increase its visibility and lead to a faster correction on Wikipedia van2011science ; wang2021science , it could also create editorial inertia if its findings are widely accepted, resulting in a slower correction.
  • Team Size. Team size is measured by counting the number of paper authors. Larger teams may contribute to greater research visibility and faster dissemination due to broader professional networks li2022untangling . However, prior studies suggest that increased team size can also complicate accountability, potentially leading to delays in retraction or correction processes mongeon2016costly .
  • Academic Domains. We include academic domains to examine whether the speed of citation correction varies across broad fields of research. Domains are derived from OpenAlex’s primary_topic field, which is based on Scopus’s ASJC classification Elsevier and assigned using OpenAlex’s topic classification methods openalexend . We focus on four high-level domains: Health Sciences, Life Sciences, Physical Sciences, and Social Sciences. To incorporate these categorical variables into our regression model, we applied one-hot encoding and omitted Health Sciences as the reference category to avoid multicollinearity. Differences across domains may reflect disparities in editorial practices, public scrutiny, or disciplinary norms. For example, Health Sciences often receive greater media and public attention, which can accelerate the detection and correction of retracted work furman2012governing . In contrast, corrections in the Social Sciences may lag due to longer publication cycles.
  • Tweet Mention Count. Tweet mention count data is collected from Altmetric LLC altmetric . To ensure methodological consistency and avoid look-ahead bias, we measured the number of tweets up to the start of each citation’s unique “at risk” period. This count serves as a proxy for the paper’s online visibility and public exposure at the moment it became problematic on Wikipedia. While a higher count of these pre-existing tweets could signal a more prominent paper that editors are more likely to scrutinize, we acknowledge that the attention generated by the retraction event itself is another important mechanism. Although modeling this second effect is a valuable direction for future work, this study focuses on the role of a paper’s pre-existing visibility. Prior research highlights the role of online engagement in scientific controversies, finding that public attention on platforms like Twitter often spikes around the time a paper is retracted peng2022dynamics . While news and blog mentions also serve as valid proxies for online engagement, we found they were highly correlated with the Tweet mention count. To avoid multicollinearity in our model, we therefore selected the Tweet mention count as the most representative measure.
3.6.2 Retraction Event Features
This feature group captures characteristics of the retraction event itself. These features provide context on the severity of the paper’s issues and the timeline of its withdrawal, which may signal to editors the urgency of a correction and thus influence its speed. Data for these features were retrieved from OpenAlex priem2022openalex and Retraction Watch oransky2012retraction .
  • Time to Retraction. Time to retraction is calculated as the time between a paper’s publication date and its official retraction date, retrieved from OpenAlex and Retraction Watch. Longer times between publication and retraction may delay the correction of citations on Wikipedia, as outdated citations may become embedded in the platform’s content. Prior studies have highlighted the negative impact of delayed retractions on scientific credibility and public trust brainard2018massive .
  • Pre-Retraction Editorial Events. Editorial events include corrections, expressions of concern, or other non-retraction updates associated with a given DOI. If such events existed, the variable was labeled as 1. Such events could serve as warnings for editors or readers, potentially influencing how quickly problematic citations are identified and addressed.
  • Retraction Reason. According to the framework presented in Zhang et al., the retraction reason can be categorized into 8 groups: “plagiarism”, “self-plagiarism”, “falsification/manipulation”, “error & unreliable results”, “authorship issues”, “ethical issues”, “others”, and “not available/lack of information” zhang2020collaboration . For this study, we focused specifically on falsification/manipulation (fraud) and error & unreliable results (error). Following this framework, we searched for the keywords “falsification,” “manipulation,” and “misconduct” in the Retraction Watch reasons to identify fraud papers, and we used the keywords “error,” “contamination,” and “unreliable” to identify error papers. These fraud and error categories were then combined into a broader category labeled “fraud & errors,” capturing retractions that may draw more editorial attention. Papers falling into this category were assigned a binary value of 1.
3.6.3 Wikipedia Page Features
This feature group captures characteristics of Wikipedia pages that may influence the speed with which retracted citations are identified and corrected. We aim to understand how Wikipedia’s collaborative and dynamic editorial processes impact the persistence of problematic citations. All features in this group are derived from Wikimedia data, retrieved through the Wikimedia REST MediaWiki and MediaWiki Action WikiAction APIs.
  • Number of Unique Editors. This variable serves as a proxy for the level of collaborative oversight on a given Wikipedia page. To ensure this was a valid baseline covariate and to avoid look-ahead bias, we measured it dynamically. For each citation pair, we calculated the number of distinct human editors who had contributed to the page up to the start of that citation’s risk period (i.e., the point after the cited paper was both retracted and present on the page). To ensure the measure reflects meaningful editorial engagement, we excluded bot and anonymous edits from this count. The rationale is that a page with a larger community of editors at the time a citation becomes at risk may experience greater collaborative oversight, potentially leading to a faster correction.
  • Controversy Noted at Initial Citation. This feature indicates whether the Wikipedia article describes the cited paper as controversial or disputed at the time it was first added to the page. We identify this by manually checking whether the initial citation frames the paper as debated, flawed, or otherwise questionable. Such editorial framing may signal early skepticism and influence the speed at which the citation is later corrected.
We excluded revision count at risk time from the model due to its near-perfect correlation with the number of unique editors at risk time.
3.7 Data Preprocessing and Model Specification
Before model fitting, we preprocessed the independent variables to ensure model stability. We applied a log-transformation (log(x+1)) to continuous variables exhibiting high positive skew, including academic citation count, team size, tweet mention count, time to retraction, and number of unique editors. Subsequently, these same continuous variables were standardized using z-score normalization to allow for the comparison of their effect sizes. Binary and dummy variables were left unchanged. Our initial sample of 845 problematic citations was reduced to a final analytical sample of 764 for our models after excluding “proactive corrections”—cases where the correction was made before the official retraction date.
Our analysis proceeds in two stages: first, we assess the suitability of the standard Cox Proportional Hazards model (PH), and second, based on diagnostic tests, we select and specify a more appropriate Accelerated Failure Time (AFT) model.
3.7.1 Model Selection
We began by fitting a Cox PH model, the semi-parametric standard for survival analysis. We then tested the model’s core assumption of proportional hazards using a statistical test based on scaled Schoenfeld residuals schoenfeld1982partial . The test revealed that our categorical variable for the academic domain exhibited non-proportional hazards (see Appendix Table A1). Specifically, the dummy variable representing the Life Sciences domain failed the test ($p=0.0347$), indicating that its effect on the hazard of correction is not constant over time.
Given this violation of a core assumption, we concluded that the Cox PH model was not the most appropriate choice for our data. We therefore turned to the family of parametric Accelerated Failure Time (AFT) models. Unlike the Cox model, which models the hazard rate, AFT models directly model the time to an event and do not rely on the proportional hazards assumption, making them a robust alternative in this context.
3.7.2 AFT Model Specification
Let $T_{i}$ denote the time (in days) between a citation $i$ entering its risk period to the first correction event on Wikipedia. To model this, we specified a general Accelerated Failure Time (AFT) model, which links the logarithm of the survival time to a set of covariates:
$\log T_{i}=\beta_{0}\;$ $+\beta_{1}\,\mathrm{OpenAccess}_{i}+\beta_{2}\,\mathrm{CitationsAtRisk}_{i}+\beta_{3}\,\mathrm{TeamSize}_{i}$
$+\beta_{4}\,\mathrm{PhysSci}_{i}+\beta_{5}\,\mathrm{LifeSci}_{i}+\beta_{6}\,\mathrm{SocSci}_{i}+\beta_{7}\,\mathrm{TweetMentions}_{i}$
$+\beta_{8}\,\mathrm{TimeToRetraction}_{i}+\beta_{9}\,\mathrm{PreRetractionEditorial}_{i}+\beta_{10}\,\mathrm{RetractionReason}_{i}$
$+\beta_{11}\,\mathrm{UniqueEditors}_{i}+\beta_{12}\,\mathrm{ControversyNoted}_{i}+\varepsilon_{i},$ (1)
where:
  • $\textit{OpenAccess}_{i}$: indicator if the paper was published open access;
  • $\textit{CitationsAtRisk}_{i}$: academic citation count of the paper at the start of the risk period;
  • $\textit{TeamSize}_{i}$: number of authors of the retracted paper;
  • $\textit{PhysSci}_{i}$, $\textit{LifeSci}_{i}$, $\textit{SocSci}_{i}$: academic domain dummies (Health Sciences as reference);
  • $\textit{TweetMentions}_{i}$: number of Twitter/X mentions at the start of the risk period;
  • $\textit{TimeToRetraction}_{i}$: days from publication to official retraction;
  • $\textit{PreRetractionEditorial}_{i}$: indicator for any editorial action prior to retraction (e.g., expression of concern);
  • $\textit{RetractionReason}_{i}$: salient retraction reason (fraud/manipulation or error/unreliable results);
  • $\textit{UniqueEditors}_{i}$: number of distinct human editors contributing before the risk period;
  • $\textit{ControversyNoted}_{i}$: indicator if the citation was framed as controversial at the time of addition;
  • $\varepsilon_{i}$: an error term following an assumed parametric distribution.
To select the most appropriate parametric distribution for our data, we fitted three distinct AFT models—Weibull, Log-Normal, and Log-Logistic—which are distinguished by the assumed distribution of the error term $\varepsilon_{i}$. We then compared their relative goodness of fit using the Akaike Information Criterion (AIC), a standard metric for model selection that penalizes model complexity akaike2003new . The Weibull AFT model yielded the lowest AIC and was therefore selected for our final analysis (see Appendix Table A2). All survival analyses were conducted in Python using the lifelines package davidson2019lifelines .
4 Results
4.1 Overview of Retracted Papers on Wikipedia
Figure 3: Share of problematic (orange) vs unproblematic (green) citations on Wikipedia.
4.1.1 Problematic Citations
We analyze the distribution of different types of retracted paper citations on Wikipedia. As shown in Figure 3, among the 1,181 citation pairs found, 845 were identified as problematic citations, with 608 cited before retraction and 237 cited after retraction without proper in-text warnings. This means that 71.6% of all collected citations represent problematic cases, where retracted papers were either cited before their retraction status became known or cited afterwards without adequate warnings to readers. 51.5% of all citations fall into the former category. These citations often persist due to a lack of systematic monitoring, leaving articles exposed to outdated or invalidated information. In contrast, 28.3% of the citations were non-problematic, meaning that their retracted status was appropriately mentioned at the time of citation. Critically, the category of citations introduced after a paper’s retraction but without proper in-text warnings constitutes 20.2% of all citations in our sample, revealing the limitations of current editorial and automated interventions in flagging and fixing such citations.
Figure 4: Comparison of the ratios of retracted and non-retracted papers cited on Wikipedia across academic domains. The retracted ratio represents the fraction of all retracted papers within a domain that are cited on Wikipedia (blue). In contrast, the non-retracted ratio is the fraction of non-retracted papers cited on Wikipedia (red).
4.1.2 Citation of Retracted vs Non-retracted Papers Across Academic Domains
We examined how retracted papers are cited on Wikipedia relative to non-retracted papers across major academic domains. We compute the retracted ratio as the proportion of retracted papers within each domain that Wikipedia cites. Similarly, the non-retracted ratio is calculated as the proportion of non-retracted papers within each domain that Wikipedia cites. For example, in Life Sciences, the retracted ratio refers to the fraction of all retracted Life Sciences papers that appear as citations on Wikipedia. This ratio is 0.0206, representing the fraction of all retracted Life Sciences papers appearing as Wikipedia citations. We used a separate dataset acquired from Altmetric for this analysis, and we collected 538 retracted papers indexed in Altmetric. Figure 4 compares ratios between retracted and non-retracted papers across four academic domains: Physical Sciences, Social Sciences, Health Sciences, and Life Sciences.
The Health Sciences domain stands out with the highest ratio of retracted papers cited on Wikipedia, at 0.0220, compared to a much lower 0.0064 ratio for non-retracted papers. This disparity suggests a heightened vulnerability of this domain to problematic citations. Since retracted Health Sciences papers are 3.44 times more likely to be cited on Wikipedia than non-retracted ones, formal retractions in this high-stakes field are not reliably reflected in Wikipedia’s citation practices.
In contrast, the Physical and Social Sciences exhibit near parity between the ratios of retracted and non-retracted papers cited (0.0064 vs. 0.0063 for Physical Sciences; 0.0093 vs. 0.0064 for Social Sciences), suggesting more balanced editorial attention. Life Sciences show a modest gap, with 0.0206 for retracted and 0.0158 for non-retracted papers, highlighting that while problematic citations are prevalent, they are not disproportionately overrepresented relative to overall citation activity in the domain.
These findings reveal that citation persistence varies by academic field, with Health Sciences exhibiting the most acute overrepresentation of retracted work. Given Wikipedia’s widespread use as a public-facing medical reference, addressing this gap is particularly crucial for maintaining information integrity.
4.2 Correcting Problematic Citations on Wikipedia
Figure 5: Distribution of corrections of problematic citations on Wikipedia.
4.2.1 Distribution of the Corrections
As shown in Figure 5, 32.9% of problematic retracted paper citations on Wikipedia ($n=278$) remained uncorrected when the data was collected, exposing readers to possible misinformation. While 12.2% are addressed through in-text retraction acknowledgments ($n=103$) and 54.9% are completely removed ($n=464$), the non-negligible percentage of unaddressed cases reveals substantial gaps in editorial intervention.
Figure 6: Kaplan-Meier survival curve of problematic citations. The y-axis shows the proportion of citations that remain uncorrected over time. The x-axis represents the time to correction in days, measured from the start of the risk period for each citation, which is defined as the later of the paper’s official retraction date or the citation’s addition to Wikipedia.
4.2.2 Descriptive Analysis of Correction Times
Our final analytical sample consists of 764 problematic citations. Of these, 486 were corrected during our observation period (the event), while 278 remained uncorrected and were right-censored at the end of the study. To visualize this process, we use the Kaplan-Meier estimator kaplan1958nonparametric , a standard non-parametric method for plotting the probability of an event not yet occurring over time while properly accounting for censored data. The resulting survival curve for our sample is shown in Figure 6.
The curve reveals a rapid rate of correction shortly after a citation’s source paper has been retracted and is present on a Wikipedia page, followed by a long tail of persistence for those citations that are not addressed quickly. The median time to correction for problematic citations was found to be 1,344 days (approximately 3.68 years). While many corrections occur swiftly, the distribution is highly skewed. The longest observed duration for a citation to remain uncorrected before being censored was 6,107 days (over 16.73 years). Notably, the Kaplan-Meier analysis estimates that the probability of a citation remaining uncorrected after 1,000 days (approximately 2.74 years) is 67.3%, indicating persistent challenges in timely intervention.
Figure 7: Coefficients from the Weibull Accelerated Failure Time (AFT) model predicting the time to correction for problematic citations (Concordance = 0.671; log-likelihood ratio test: $p<0.001$). Feature categories group coefficients. Bars indicate the coefficient estimates; error bars represent a $95\%$ confidence interval. The shapes and colors of the data points denote feature groups. Features with negative coefficients are associated with an accelerated time to correction (i.e., faster correction), while positive coefficients indicate a decelerated time (i.e., slower correction).
4.2.3 Factors Influencing the Time to Correction: AFT Model Results
To identify factors that predict the time until a problematic citation is corrected, we fitted an AFT model with an assumed Weibull distribution for the survival times, which directly models the speed of correction. The results are interpreted as time ratios, where a ratio below 1 indicates a faster correction and a ratio above 1 indicates a slower correction.
The overall model is statistically significant ($\text{log-likelihood ratio test}=117.551$, $p<0.001$) and demonstrates moderate predictive accuracy with a concordance index of 0.671. While some independent variables are modestly correlated (see Appendix Figure B1), all variance inflation factors (VIFs) are well below the standard threshold of 5 (see Appendix Table B1), indicating that multicollinearity is not a concern. The following sections detail the effects of specific paper-, retraction event-, and Wikipedia page-level features on the time to correction. Figure 7 presents the results, which are discussed below, broken down by feature groups.
Paper Features
The visibility and accessibility of a paper are significantly associated with its correction speed. Papers that are open access are corrected significantly faster than those behind a paywall ($\text{Time Ratio}=0.522$, $p<0.001$). Similarly, a higher level of Twitter/X attention prior to retraction is associated with a faster correction time ($\text{Time Ratio}=0.567$, $p<0.001$), suggesting that public attention on social media may serve as an effective signal to Wikipedia editors.
The academic domain of a paper also shapes the speed of its correction. Citations of papers in the Physical Sciences ($\text{Time Ratio}=0.359$, $p<0.001$) and Social Sciences ($\text{Time Ratio}=0.525$, $p<0.05$) are corrected significantly faster relative to Health Sciences, while citations in the Life Sciences ($\text{Time Ratio}=0.939$, $p=0.769$) show no significant difference with respect to Health Sciences. These findings suggest that domains with greater online or editorial attention may experience more timely corrections, while others may be overlooked. As our descriptive analysis in Figure 4 shows, the Health Sciences domain exhibits the most acute over-representation of retracted work, and we selected it as the reference category for our model. All reported domain effects are therefore relative to this baseline of heightened vulnerability.
Conversely, a higher academic citation count at the start of the risk period (i.e., the point after a paper is both retracted and cited on Wikipedia) is linked to a slower correction ($\text{Time Ratio}=1.352$, $p<0.05$). This may indicate that more highly-cited papers are more deeply embedded in Wikipedia’s content, making editors more hesitant to remove or amend them. The research team size was not found to have a significant effect on the correction time.
Retraction Event Features
Not all retractions are treated equally on Wikipedia. Papers retracted for reasons such as fraud, manipulation, or unreliable results are corrected significantly faster ($\text{Time Ratio}=0.786$, $\textit{p}<0.05$), suggesting that the severity of the retraction notice influences editorial response. When retraction notices explicitly flag misconduct or data concerns, editors may be more likely to notice and address the problematic citation, likely due to heightened visibility or urgency. However, this does not diminish the importance of addressing other types of retractions, such as those resulting from procedural or ethical issues, which may take longer to be recognized or corrected.
Neither the time between a paper’s publication and retraction ($\text{Time Ratio}=0.854$, $p=0.186$) nor the presence of a pre-retraction editorial event ($\text{Time Ratio}=0.899$, $p=0.707$) had a significant effect on the time to correction.
Wikipedia Page Features
The characteristics of the Wikipedia page hosting the citation are significant predictors of correction speed. A greater number of unique human editors contributing to a Wikipedia page is significantly associated with a faster correction ($\text{Time Ratio}=0.789$, $p<0.01$). This suggests that broader editorial engagement may enhance collective oversight, thereby increasing the likelihood that problematic citations are identified and addressed more promptly.
Furthermore, we find that citations that were flagged as controversial, e.g., those accompanied by dispute tags or appearing in debated content, are corrected significantly faster ($\text{Time Ratio}=0.498$, $p<0.05$). This suggests that early signals of conflict or scrutiny may increase the visibility of problematic content and prompt corrective action. Together, these findings highlight that the scale of human participation and oversight is a key factor associated with the accelerated repair of retracted citations on Wikipedia.
5 Discussion
The persistence of retracted paper citations on Wikipedia represents a critical vulnerability in the community’s ability to maintain epistemic integrity collaboratively. While Wikipedia is often recognized for its agility in responding to breaking news and high-salience events keegan2012staying ; twyman2017blm , our findings reveal a much slower and uneven response to post-publication knowledge failures. Specifically, 71.6% of the retracted paper citations we examined were problematic. Our survival analysis shows the extent of this delay: the median time to correction is 3.68 years, and a “long tail” of persistence indicates that 67.3% of problematic citations remain uncorrected after nearly 2.74 years. These long delays reflect a gap between Wikipedia’s technical infrastructure for correction and the community’s ability to mobilize editors. Given that Wikipedia continues to serve as a key access point for scientific information, this fragility in information repair raises essential concerns for both end users and the broader knowledge ecosystem.
Our findings can be understood through the lens of a classic coordination challenge in large-scale collaborative systems, where achieving quality is fundamentally dependent on successful coordination between contributors kittur2008harnessing . This is more than a simple information routing issue, where public retraction notices fail to reach the right editors and articles ackerman2013sharing . The challenge is also one of collective sensemaking and action keegan2013hot ; weick2005organizing . Even when editors are aware of a retraction, they may disagree on the appropriate response—whether to remove the citation entirely or add a warning viegas2004studying ; kittur2008harnessing —or on the urgency of the repair compared to other maintenance tasks, as volunteer attention is a scarce resource in large-scale peer production systems kittur2007he . Therefore, the persistence of these citations reveals a breakdown not only in information flow, but also in the community’s ability to achieve consensus and prioritize corrective action. The remainder of this discussion interprets our empirical results through this framework, examining how different factors related to the paper, the retraction event, and the community enhance or inhibit this coordination.
Our analysis reveals that the prevalence of problematic retracted citations is not evenly distributed across academic domains. The Health Sciences domain exhibits the highest rate of retracted paper citations on Wikipedia relative to the total number of retracted papers in that field. In fact, retracted Health Science papers are over three times as likely to be cited on Wikipedia as non-retracted ones. This overrepresentation is particularly troubling given Wikipedia’s prominent role as a public health information source heilman2011wikipedia . While our results show that citations in the Physical and Social Sciences are corrected more quickly, the elevated citation rate in the Health Sciences still increases overall exposure to retracted findings. These findings underscore the necessity for domain-specific strategies to mitigate citation decay and enhance epistemic integrity in high-risk areas, such as health communication.
We also find that not all retractions are treated equally. Citations to papers retracted for high-salience reasons, such as fraud, data manipulation, or serious methodological error, are corrected significantly faster than those withdrawn for ambiguous or technical issues, like authorship disputes or duplication. This suggests that Wikipedia editors may prioritize fixing citations to papers that seem more serious or clearly problematic, such as those involving fraud or significant errors, while not prioritizing less severe or technical retractions. This editing pattern reflects a common trend in collaborative work during crises, where people focus more on problems that seem urgent or clearly defined, and are less likely to act when issues are more complex to recognize or explain twyman2017blm ; keegan2013hot .
A particularly robust finding is the role of human editorial activity. Articles edited by more unique contributors have their retracted citations corrected more quickly. This reinforces established CSCW literature on the benefits of having a larger number of active editors working on an article, which enhances content quality through diversity and redundancy halfaker2013rise ; forte2012coordination . However, it also highlights a structural inequality: many Wikipedia articles receive little editorial attention and thus lack the oversight necessary for timely citation repair. These gaps in attention create a long tail of vulnerability, where outdated or invalid knowledge objects remain visible and unchallenged.
A key consideration is the role of automated tools, such as RetractionBot Headbomb2024Signpost . This bot exemplifies the specialized roles that automated agents play in Wikipedia’s quality control ecosystem zheng2019roles . It primarily serves an editorial audience. By systematically adding a prominent template to the reference section, the bot is highly effective at its specific task of signaling a source’s retracted status to editors engaged in verification and maintenance. The promise of such tools lies in this efficiency for well-defined tasks, and the community has become highly reliant on them for routine maintenance. For instance, the temporary failure of another specialized bot, an anti-vandalism tool, was shown to nearly double the time it took for the Wikipedia community to revert malicious edits geiger2013levee . However, our work highlights a persistent gap between the effectiveness of automation for these specific, often editor-facing tasks and the challenges of repairing more nuanced, epistemic issues for a general reader. This distinction is key: while a bot can efficiently apply a “technical flag,” this action is distinct from the substantive, contextual repair required to update an article’s main text. This reveals a fundamental sociotechnical challenge in designing workflows that successfully translate automated alerts into the complex, collaborative human judgment needed for complete epistemic repair.
Our analysis also points to the role of external visibility cues in prompting correction. Citations to papers that received significant social media attention—particularly on Twitter/X—were corrected more quickly. This finding suggests that social signals external to Wikipedia can help surface retractions to editors and may serve as an informal triage mechanism. Similarly, our finding that open access papers are corrected faster implies that free access can accelerate editorial verification and repair. These findings suggest that social signals and open accessibility can help surface retractions to editors. However, not all forms of visibility lead to faster corrections. We found that papers with a higher academic citation count are actually corrected slower, perhaps because their claims are more accepted and editors are more hesitant to challenge them. These patterns emphasize the importance of salience, not just in the nature of the retraction but in its discoverability.
Together, these findings highlight a multilayered vulnerability in Wikipedia’s citation infrastructure—one rooted not in neglect, but in the subtle interplay between visibility, editorial capacity, and the nature of retraction events. Ultimately, this challenge is not unique to Wikipedia. Similar coordination issues likely exist in other distributed knowledge systems, such as open-source software projects managing bug fixes across dependencies, or collaborative fact-checking platforms striving to correct misinformation that has already propagated widely forte2012coordination ; kittur2008harnessing . Future work should continue to explore how to better surface and prioritize these risks, especially in domains where epistemic failure carries real-world consequences. Our work thus uses Wikipedia as a case study to illuminate a general vulnerability in the maintenance of large-scale, collaboratively produced knowledge, highlighting the critical need for systems that not only detect errors but also facilitate the complex social coordination required to repair them ackerman2013sharing ; jackson201411 .
5.1 Design Implications
Our findings suggest several design opportunities to address a fundamental coordination challenge of how to efficiently identify and repair citations to retracted scientific papers within Wikipedia’s maintenance ecosystem. The core issue is not a lack of information—retraction notices are public and trackable—but a two-fold coordination challenge: first, a failure in connecting this critical information with the specific contexts and contributors responsible for repair; second, the subsequent challenge of achieving collective sensemaking to agree upon and prioritize a corrective action. Our survival analysis quantifies the consequences of this breakdown, revealing that correction delays can last for years. Crucially, our model demonstrates that signals of human attention may bridge the persistent gap between information availability and coordinated action. The following design directions are therefore aimed at strengthening the sociotechnical mechanisms that facilitate this crucial routing and coordination work.
5.1.1 Improving Retraction Visibility in Context
Our analysis reveals that early, human-generated warnings are effective; citations flagged as controversial are corrected significantly faster. However, such in-text flags are applied inconsistently. This suggests a need for tools that make these contextual warnings easier to create and more visible. To solve this information routing problem, Wikipedia’s editing interface could surface retraction indicators directly within citation markup or during editing interactions. For example, editors could receive contextual alerts (e.g., a badge or tooltip) when adding or modifying citations to retracted sources, along with links to the official retraction notice. By embedding retraction cues within the editing workflow, these changes would align with CSCW insights about situated action and the importance of interface-level nudges for distributed maintenance work geiger2017operationalizing .
5.1.2 Prioritizing Correction Work through Dashboard Support
Our analysis shows that pages edited by a larger number of unique human contributors have their retracted citations corrected more quickly, reinforcing CSCW research on the benefits of crowd diversity and sustained participation in maintaining information quality forte2012coordination ; halfaker2013rise . However, retracted citations are not evenly distributed: some academic domains—most notably Health Sciences—exhibit disproportionately high rates of retracted paper citations relative to non-retracted ones. While our AFT model shows citations in other domains are corrected faster, their elevated citation rate increases overall exposure to problematic knowledge, particularly in the Health and Life Sciences, which our model found to have similarly slow correction speeds.
These patterns suggest that editorial attention is both influential and unevenly distributed. Wikipedia could address this by integrating retraction data into community coordination tools, such as dashboards or watchlists, to help surface under-attended articles that cite retracted work. To help editors triage their efforts, pages could be prioritized using a risk score model. This model should combine standard metrics, such as article traffic, with the novel predictors identified in our analysis. For instance, our findings suggest that the highest priority should be given to citations of papers with a high academic citation count and paywalled access, especially when they appear in articles with a low number of unique editors, as our AFT model shows these factors are significantly associated with slower correction times. This approach reflects successful CSCW strategies for managing maintenance workloads and prioritizing collaborative repair tasks across distributed systems teblunthuis2018revisiting .
5.1.3 Automating Credibility Cues through Structured Metadata
While editors can manually apply retraction templates, there is no consistent or automated mechanism to propagate retraction warnings across articles that cite the same paper. Our study reveals that corrections are uneven and often delayed, sometimes by years, resulting in gaps in credibility signaling. To reduce this burden, citation templates could be expanded to automatically ingest and display structured metadata from Crossref, Retraction Watch, or similar sources. If a DOI is retracted, the template could render a default warning, either inline or as a footnote, linking to the retraction notice. This would both standardize correction practices and reduce the cognitive load for editors.
5.1.4 Surfacing External Signals for Epistemic Risk
Finally, we find that external signals of visibility and accessibility are linked to correction speed. Retracted papers mentioned more frequently on Twitter/X are corrected significantly faster, while paywalled papers are corrected more slowly. This suggests that social media attention functions as an external signal, surfacing now-invalidated citations, while a lack of open access may act as a barrier to verification and repair. Future tools might incorporate external alerting systems, integrating tweet activity and public discussions on platforms such as blogs or PubPeer regarding specific retractions into internal Wikipedia tools. This could help editors recognize emerging credibility issues even in articles that lack sustained editorial engagement.
Wikipedia’s infrastructure does not lack data about retractions—it lacks mechanisms to surface, route, and act on that data within existing workflows. By embedding signals into interfaces, prioritizing repair work through community-facing tools like dashboards and watchlists, and linking Wikipedia’s knowledge infrastructure with external retraction metadata, designers and editors can enhance the platform’s capacity for timely and collaborative epistemic repair.
5.2 Limitations and Future Work
While comprehensive in its analysis of the persistence and correction of retracted paper citations on Wikipedia, this study has several limitations that warrant further investigation.
First, our dataset is based on integrations across Retraction Watch, Crossref Event Data, Altmetric, and OpenAlex. While these sources provide broad coverage of retraction events and citation metadata, they are not exhaustive. In particular, both Crossref and Altmetric primarily track DOI-based citations, which means that retracted papers without DOIs, such as older publications or those in certain disciplines, are likely underrepresented in our analysis. Similarly, Wikipedia citations that refer to retracted papers through PubMed IDs, title strings, or bare URLs may be missed if they are not linked through standard identifiers. Future work could expand coverage by incorporating additional retraction databases (e.g., PubMed retraction tags), parsing free-text citation formats, or combining structured detection with manual validation to better capture edge cases.
Second, although our modeling identifies significant predictors of the speed with which retracted citations are corrected, it cannot fully account for the editorial processes and sociotechnical dynamics that shape citation repair on Wikipedia. Decisions to add, retain, or remove citations are often entangled with disputes over content reliability, page notability, or conflicting interpretations of sourcing policies. Our analysis does not surface these nuanced deliberations, which may vary across topics, editor groups, or article histories. Qualitative research—such as interviews with active editors, analysis of talk page discussions, or observational studies of how citation templates are used—could provide deeper insight into how retraction cues are interpreted and acted upon in practice.
Another limitation lies in the scope of our survival analysis model (specifically, the Weibull AFT model), which, while effective for identifying associations, does not capture the complex interactions between features. For instance, factors such as citation timing, editorial activity, and retraction reasons may interact in non-linear ways. Future work could employ more sophisticated machine-learning approaches or mixed-methods frameworks to uncover deeper patterns and relationships in the data.
By addressing these methodological limitations, future research can develop a more holistic understanding of the complex interplay between individual actions, community norms, and platform design in the ongoing process of collaborative knowledge maintenance.
6 Conclusion
This study examined the presence and persistence of retracted paper citations on Wikipedia, identifying key factors influencing the speed of corrective actions. Our findings reveal that while some citations are corrected promptly, many persist uncorrected for extended periods, with a median time to correction of over 3.68 years (1,344 days). The speed of correction is significantly accelerated by factors such as a salient retraction reason, high Twitter/X attention, open access publication, a greater number of unique Wikipedia editors, and whether the paper was initially noted as controversial. Conversely, a high academic citation count is associated with a slower correction time. These findings highlight a fundamental disconnect between the public availability of retraction data and the community’s collaborative workflows, which often fail to translate this information into the reader-facing edits required for a complete repair. Addressing this challenge requires moving beyond simple error flagging towards a tighter integration of information systems and human editorial processes. Design interventions that surface different types of signals to support editorial work—such as contextual signals embedded in editing interfaces, structured data signals within citation templates, and prioritized attention signals on community dashboards—could significantly improve responsiveness to invalidated science. By combining better information routing with sustained human oversight, Wikipedia can more effectively manage citation-level credibility failures and strengthen its role as a reliable public knowledge platform.
Acknowledgments
We are deeply grateful to Aaron Shaw for insightful discussions and suggestions. We also thank the anonymous reviewers at ICWSM and IC2S2, as well as the research talk participants at IC2S2, for their constructive comments that helped improve this manuscript.
Appendix A Survival Analysis Model Selection Details
Table A1: Test of the Proportional Hazards Assumption. The test is based on Schoenfeld residuals. A significant p-value (p < 0.05) indicates that a variable violates the assumption.
Variable Chi-Square ($\chi^{2}$) df p-value
Open Access 0.18 1 0.67
Academic Citation Count (at risk start) 0.02 1 0.88
Team Size 1.88 1 0.17
Physical Sciences (domain) 0.47 1 0.49
Life Sciences (domain) 4.46 1 0.03
Social Sciences (domain) 0.24 1 0.62
Tweet Mention Count 1.10 1 0.29
Time to Retraction (days) 0.73 1 0.39
Pre-Retraction Editorial Event 1.01 1 0.32
Retraction Reason (salient) 1.21 1 0.27
Unique Editors (at risk start) 2.37 1 0.12
Controversy Noted Initially 0.24 1 0.63
Table A2: Comparison of Accelerated Failure Time (AFT) Model Fit. The model with the lowest Akaike Information Criterion (AIC) was selected for the final analysis.
Model Distribution AIC Concordance Log-Likelihood Ratio Test ($\chi^{2}$)
Weibull 8209.512 0.671 117.551
Log-Logistic 8310.896 0.676 134.926
Log-Normal 8478.694 0.673 130.386
Appendix B Feature Correlation and Multicollinearity
Figure B1: Correlation matrix of independent variables used in the Weibull AFT model. Most features exhibit low to moderate correlation, with no evidence of severe multicollinearity.
Table B1: Variance inflation factors (VIFs) for all independent variables. All VIFs are well below the common threshold of 5, indicating no multicollinearity.
Feature VIF
Open Access 1.09
Academic Citation Count (at risk start) 1.93
Team Size 1.34
Physical Sciences (domain) 1.27
Life Sciences (domain) 1.47
Social Sciences (domain) 1.15
Time to Retraction (days) 1.81
Pre-Retraction Editorial Event 1.24
Retraction Reason (salient) 1.20
Tweet Mention Count 1.49
Unique Editors (at risk start) 1.19
Controversy Noted Initially 1.05